# The Invisible Leash? Why RLVR May or May Not Escape Its Origin

### v3

@@ 5 Conclusion
We reveal that the current RLVR improves precision by sharpening distributions around known high-reward trajectories, yet largely preserves the base model’s support. Importantly, we found that this sharpening does not merely prune incorrect outputs—it can also concentrate probability mass on a narrower subset of correct solutions, occasionally excluding valid alternatives that the more diverse base model could still recover. Meanwhile, the divergence between token-level uncertainty and answer-level diversity indicates that stepwise stochasticity alone is insufficient for global exploration, motivating future work to bridge this gap. We suggest that to expand reasoning capabilities beyond the base model’s scope, RLVR must be coupled with explicit exploration strategies or off-policy mechanisms that seed probability mass into underrepresented regions of the solution space.
Ethics Statement
This research was conducted in accordance with established academic ethical standards using publicly available models and appropriate computational resources. All models were accessed through proper channels with necessary permissions, and outputs were analyzed solely for research purposes to understand fundamental properties of reinforcement learning techniques. We acknowledge the significant computational resources required for LLM evaluation and made efforts to optimize experimental design while maintaining scientific rigor. This work investigates limitations of current RL approaches in AI systems, and we believe transparent reporting of these constraints is essential for guiding effective future research directions, helping practitioners set appropriate expectations, and contributing to broader scientific understanding of AI capabilities. While our analysis focuses on specific model families and current benchmarks, which may limit generalizability, we encourage continued research into these limitations and the development of methods that can genuinely expand model reasoning capabilities while maintaining the beneficial precision improvements that RLVR provides.
Acknowledgment
This work was supported in part by RS-2024-00457882, National AI Research Lab Project.
@@ Appendix A Detailed Statistics for Support Dynamics
We provide full per-model statistics that underlie the aggregate values in Table 1. For each model and domain (Math, Non-Math, and Overall), we report the raw counts of correct completions across the four empirical support categories: Preservation (P), Expansion (E), Shrinkage (S), and Out-of-Support (O). From these counts, we compute the derived metrics: Support Retention Rate (SRR), Net Discovery Rate (NDR), Support Dynamic Score (SDS), and Net Support Change Rate (NSCR).
@@
Math Benchmarks 0.960 0.0037 0.0073 -0.036 1355 5 56 131
Non-Math Benchmarks 0.907 0.0288 0.0558 -0.064 1045 31 107 674
@@
Math Benchmarks 0.994 0.003 0.007 -0.003 1431 5 9 102
Non-Math Benchmarks 0.965 0.018 0.035 -0.018 1284 23 47 503
@@
Math Benchmarks 0.976 0.001 0.003 -0.023 1406 2 34 105
Non-Math Benchmarks 0.961 0.018 0.036 -0.021 1279 24 52 502
@@
Math Benchmarks 0.990 0.0035 0.0070 -0.0069 1425 5 15 102
Non-Math Benchmarks 0.992 0.0030 0.0060 -0.0050 993 3 8 399
@@
Math Benchmarks 0.992 0.0057 0.0112 -0.0028 1407 8 12 120
Non-Math Benchmarks 0.990 0.0074 0.0148 -0.0028 1067 8 11 317
@@
Math Benchmarks 0.887 0.110 0.166 -0.072 761 83 104 599
@@
Dataset Step Pass@256 Support Dynamics Metrics Support Counts
RLVR SRR NDR SDS NSCR P E S O
AIME24 0 83.33 - - - - - - - -
30 80.00 0.960 0.000 0.000 -0.040 24 0 1 5
60 80.00 0.960 0.000 0.000 -0.040 24 0 1 5
90 83.33 1.000 0.000 0.000 0.000 25 0 0 5
120 80.00 0.960 0.000 0.000 -0.040 24 0 1 5
150 80.00 0.960 0.000 0.000 -0.040 24 0 1 5
180 83.33 1.000 0.000 0.000 0.000 25 0 0 5
210 76.67 0.920 0.000 0.000 -0.080 23 0 2 5
240 73.33 0.880 0.000 0.000 -0.120 22 0 3 5
270 73.33 0.880 0.000 0.000 -0.120 22 0 3 5
300 73.33 0.880 0.000 0.000 -0.120 22 0 3 5
AIME25 0 70.00 - - - - - - - -
30 70.00 0.905 0.095 0.172 0.000 19 2 2 7
60 73.33 0.905 0.136 0.237 0.042 19 3 2 6
90 70.00 0.905 0.095 0.172 0.000 19 2 2 7
120 73.33 0.905 0.136 0.237 0.042 19 3 2 6
150 73.33 0.905 0.136 0.237 0.042 19 3 2 6
180 63.33 0.905 0.000 0.000 -0.095 19 0 2 9
210 66.67 0.905 0.050 0.095 -0.045 19 1 2 8
240 66.67 0.905 0.050 0.095 -0.045 19 1 2 8
270 63.33 0.905 0.000 0.000 -0.095 19 0 2 9
300 66.67 0.864 0.095 0.172 -0.042 18 2 3 7
AMC23 0 97.50 - - - - - - - -
30 97.50 1.000 0.000 0.000 0.000 39 0 0 1
60 97.50 1.000 0.000 0.000 0.000 39 0 0 1
90 97.50 1.000 0.000 0.000 0.000 39 0 0 1
120 97.50 1.000 0.000 0.000 0.000 39 0 0 1
150 97.50 1.000 0.000 0.000 0.000 39 0 0 1
180 97.50 1.000 0.000 0.000 0.000 39 0 0 1
210 97.50 1.000 0.000 0.000 0.000 39 0 0 1
240 97.50 1.000 0.000 0.000 0.000 39 0 0 1
270 97.50 1.000 0.000 0.000 0.000 39 0 0 1
300 97.50 1.000 0.000 0.000 0.000 39 0 0 1
MATH500 0 99.20 - - - - - - - -
30 98.80 0.996 0.000 0.000 -0.004 494 0 2 4
60 98.80 0.996 0.000 0.000 -0.004 494 0 2 4
90 99.20 0.998 0.002 0.004 0.000 495 1 1 3
120 99.00 0.996 0.002 0.004 -0.002 494 1 2 3
150 98.80 0.996 0.000 0.000 -0.004 494 0 2 4
180 99.00 0.998 0.000 0.000 -0.002 495 0 1 4
210 98.80 0.994 0.002 0.004 -0.004 493 1 3 3
240 98.60 0.992 0.002 0.004 -0.006 492 1 4 3
270 98.40 0.992 0.000 0.000 -0.008 492 0 4 4
300 97.80 0.984 0.002 0.004 -0.014 488 1 8 3
Minerva 0 62.50 - - - - - - - -
30 62.13 0.935 0.059 0.111 -0.006 159 10 11 92
60 59.93 0.929 0.031 0.059 -0.040 158 5 12 97
90 61.76 0.924 0.065 0.122 -0.011 157 11 13 91
120 62.87 0.953 0.053 0.100 0.006 162 9 8 93
150 62.50 0.947 0.053 0.100 0.000 161 9 9 93
180 59.56 0.906 0.049 0.094 -0.045 154 8 16 94
210 60.29 0.924 0.043 0.082 -0.034 157 7 13 95
240 61.40 0.941 0.042 0.080 -0.017 160 7 10 95
270 59.19 0.906 0.043 0.083 -0.051 154 7 16 95
300 58.46 0.912 0.025 0.049 -0.063 155 4 15 98
Olympiad 0 88.00 - - - - - - - -
30 86.67 0.963 0.022 0.043 -0.015 572 13 22 68
60 86.81 0.960 0.027 0.053 -0.013 570 16 24 65
90 86.07 0.961 0.017 0.034 -0.022 571 10 23 71
120 86.52 0.961 0.022 0.044 -0.016 571 13 23 68
150 85.48 0.955 0.017 0.034 -0.028 567 10 27 71
180 84.00 0.944 0.011 0.021 -0.045 561 6 33 75
210 82.81 0.926 0.016 0.032 -0.058 550 9 44 72
240 82.81 0.931 0.011 0.021 -0.058 553 6 41 75
270 81.33 0.912 0.013 0.025 -0.075 542 7 52 74
300 79.70 0.897 0.009 0.018 -0.093 533 5 61 76
Appendix B Experimental Details
B.1 Evaluation Settings
B.2 Datasets
Math benchmarks.
Non-math benchmarks.
Reasoning Gym.
For Reasoning Gym (Stojanovski et al., 2025), we employ the easy set from the version updated after commit 17a8431 in its repository as our default task configuration. This choice ensures consistency with the default task configuration used in prior evaluations, maintaining comparable experimental conditions. Additionally, we utilize the hard set as our challenging evaluation benchmark for further evaluations.
B.3 Answer Processing Enhancement in Reasoning Gym
We identified significant evaluation challenges when testing the base model on Reasoning Gym. The ProRL model, having been trained on Reasoning Gym data, predominantly produces responses that conform to the expected format, leading to much higher accuracy scores. In contrast, the base model struggled with format adherence due to insufficiently detailed prompts, and its limited 1.5B parameter capacity made it particularly susceptible to evaluation inconsistencies. To address these issues, we enhanced both the answer extraction protocol and prompt design to ensure fair and objective accuracy assessments across both models. This causes the differences of ProRL’s reported performance and our evaluation results in Reasoning Gym.
B.3.1 General Answer Extraction Protocol
[…]
As a final fallback, the system extracts content following </think> tags when all previous methods fail and the response contains these markers. This safety mechanism captures base model responses that ignore formatting requirements in lengthy tasks.
B.3.2 Task-Specific Processing Modifications
Our core answer processing pipeline applies to both models, with additional processing steps designed primarily to address format compatibility issues commonly encountered with base model responses. Specifically, the processing logic for each task is enhanced as follows:
dicedice
prime_factorizationprime\_factorization
palindrome_generationpalindrome\_generation
The ground truth format expects palindromic character strings (sequences that read the same forwards and backwards). We remove excess whitespace to address frequent spacing issues in base model responses. This transformation converts spaced responses like “k h g a g h k” to “khgaghk”, preventing string reversibility judgment failures that occur when spaces interfere with palindrome verification.
advanced_geometryadvanced\_geometry
[…]
power_functionpower\_function
The ground truth format uses e-notation scientific notation. We convert mixed and arithmetic symbol scientific notation to ensure format consistency. The system transforms patterns like $2.36\times 10^{-16}$” or $1.5\times 10^{5}$” to e-notation format (“-2.36e-16”, “1.5e5”), preventing numerically correct but format-incompatible evaluation errors when base models use mixed and arithmetic symbols for scientific notation.
arc_1darc\_1d
[…]
boxnetboxnet
The ground truth format requires dictionary list formatting [{key: value}, ...]. We implement comprehensive JSON format cleaning to meet these evaluation requirements. Our processing pipeline includes several steps: rejecting pure numerical responses to prevent non-JSON format interference; removing JSON markdown wrappers that eliminate ‘‘‘json {content} ‘‘‘ markers; converting single dictionaries to single-element dictionary lists (dict $\rightarrow$ [dict]); and filtering illegal elements by removing non-dictionary components from JSON lists. Additionally, we clean nested structure values within individual dictionary entries. For nested lists, we extract the first element as the value ([{key1: [value1, value2, ...]}, ...] $\rightarrow$ [{key1: value1}, ...]). For nested dictionaries, we select matching key values when available ([{key1: {key1: value1, key2: value2, ...}}, ...] $\rightarrow$ [{key1: value1}, ...]) or default to the first element value when keys don’t match ([{key1: {key2: value2, key3: value3}}, ...] $\rightarrow$ [{key1: value2}, ...]). These modifications preserve model response content to the maximum extent while ensuring ground truth format compliance.
B.4 Entropy Analysis
Setup.
Models.
We evaluate a diverse set of reasoning models to understand the entropy characteristics across different training paradigms and parameter scales, as summarized in the following table.
Table 14: Models evaluated in the entropy analysis.
Name Full Model Name Type Parameters
DeepSeek-1.5B DeepSeek-R1-Distill-Qwen-1.5B Base 1.5B
ProRL-1.5B Nemotron-Research-Reasoning-Qwen-1.5B RLVR 1.5B
DeepSeek-7B DeepSeek-R1-Distill-Qwen-7B Base 7B
AceReason-7B AceReason-Nemotron-7B RLVR 7B
Skywork-OR1-7B Skywork-OR1-7B RLVR 7B
DeepSeek-14B DeepSeek-R1-Distill-Qwen-14B Base 14B
AceReason-14B AceReason-Nemotron-14B RLVR 14B
Qwen2.5-32B Qwen2.5-32B Base 32B
DAPO-32B DAPO-Qwen-32B RLVR 32B
Entropy Computation.
For token-level entropy computation, we employ teacher-forcing to obtain probability estimates. Specifically, after generating the 32 completions with the specified sampling parameters, we feed each generated sequence back to the model and perform a single forward pass to compute the probability distribution over the vocabulary at each token position. Answer-level entropy is computed by first extracting the final answer from each completion using Lighteval (Habib et al., 2023), then calculating the entropy over the distribution of unique answers across the 32 completions. This approach allows us to compute both token-level and answer-level entropy directly from the model’s probability distributions without introducing additional sampling variance.
@@ C.1 Support Preservation: Why RLVR Rarely Discovers New Modes
By construction we initialize the RLVR policy to the base model as $\pi_{\theta_{0}}(y\mid x)\;=\;q(y\mid x)$. Hence
@@
This manifests as a trade-off between sharpness and diversity: RLVR can improve pass@1 by concentrating mass on known high-reward modes but tends to reduce pass@k performance for larger $k$, where broader coverage is beneficial. By contrast, the base model may occasionally sample correct answers from its long-tail distribution, giving it a statistical edge under high-$k$ evaluations (Yue et al., 2025a; Liu et al., 2025). This asymptotic upper bound captures a ceiling: no matter how many samples are drawn, the RLVR-trained model cannot exceed the base model’s pass@k in the limit.
@@ C.4 Estimating the Sampling Threshold $\epsilon$ from pass@k
Appendix D The use of LLMs
This study utilized large language models solely for providing assistance with minor language enhancements. All content has undergone human review, verification, and further modification.

### v4

@@ 6 Conclusion
We show that current RLVR increases precision by sharpening probability mass around known high-reward trajectories while largely preserving the base model’s support. Crucially, this sharpening goes beyond pruning incorrect outputs: it concentrates mass on a narrower subset of correct solutions, sometimes excluding valid alternatives that the more diverse base model could recover. Moreover, the gap between token-level uncertainty and answer-level diversity suggests that stepwise stochasticity alone is insufficient for global exploration. To extend reasoning beyond the base model’s scope, RLVR must be paired with explicit exploration strategies or off-policy mechanisms that allocate probability mass to underrepresented regions of the solution space.
Impact Statements
This work examines the strengths and limitations of reinforcement learning with verifiable rewards (RLVR) in large language models, focusing on how such training reshapes the accessibility of correct solutions rather than introducing new capabilities. By providing empirical evidence that RLVR primarily improves precision while potentially narrowing solution diversity, our findings aim to help researchers and practitioners better understand the trade-offs involved in deploying RLVR-based systems. We do not anticipate direct negative societal impacts from this analysis; instead, we hope it encourages more transparent evaluation practices and motivates the development of training methods that balance reliability with broader reasoning coverage.
@@ Appendix A Detailed Statistics for Support Dynamics
We provide full per-model statistics that underlie the aggregate values in Tab. 1. For each model and domain (Math, Non-Math, and Overall), we report the raw counts of correct completions across the four empirical support categories: Preservation (P), Expansion (E), Shrinkage (S), and Out-of-Support (O). From these counts, we compute the derived metrics: Support Retention Rate (SRR), Net Discovery Rate (NDR), Support Dynamic Score (SDS), and Net Support Change Rate (NSCR).
@@
Math 0.960 0.0037 0.0073 -0.036 1355 5 56 131
Non-Math 0.907 0.0288 0.0558 -0.064 1045 31 107 674
@@
Math 0.994 0.003 0.007 -0.003 1431 5 9 102
Non-Math 0.965 0.018 0.035 -0.018 1284 23 47 503
@@
Math 0.976 0.001 0.003 -0.023 1406 2 34 105
Non-Math 0.961 0.018 0.036 -0.021 1279 24 52 502
@@
Math 0.990 0.0035 0.0070 -0.0069 1425 5 15 102
Non-Math 0.992 0.0030 0.0060 -0.0050 993 3 8 399
@@
Math 0.992 0.0057 0.0112 -0.0028 1407 8 12 120
Non-Math 0.990 0.0074 0.0148 -0.0028 1067 8 11 317
@@
Math 0.880 0.098 0.177 -0.022 761 83 104 599
Appendix B Experimental Details
B.1 Evaluation Settings
B.2 Datasets
Math benchmarks.
Non-math benchmarks.
Reasoning Gym.
For Reasoning Gym (Stojanovski et al., 2025), we employ the easy set from the version updated after commit 17a8431 in its repository as our default task configuration. This choice ensures consistency with the default task configuration used in prior evaluations, maintaining comparable experimental conditions. Additionally, we use the hard set as our challenging evaluation benchmark.
B.3 Answer Processing Enhancement in Reasoning Gym
We identified significant evaluation challenges when testing the base model on Reasoning Gym. The ProRL model, trained on Reasoning Gym data, predominantly produces responses that conform to the expected format, resulting in much higher accuracy. In contrast, the base model struggled with adherence to the format due to insufficiently detailed prompts, and its limited 1.5B parameter capacity made it particularly susceptible to evaluation inconsistencies. To address these issues, we enhanced both the answer extraction protocol and prompt design to ensure fair and objective accuracy assessments across both models. This causes the differences of ProRL’s reported performance and our evaluation results in Reasoning Gym.
B.3.1 General Answer Extraction Protocol
[…]
As a final fallback, the system extracts content following </think> tags when all previous methods fail, and the response contains these markers. This safety mechanism captures base model responses that ignore formatting requirements in lengthy tasks.
B.3.2 Task-Specific Processing Modifications
Our core answer-processing pipeline applies to both models, with additional steps designed primarily to address format compatibility issues commonly encountered in base-model responses. Specifically, the processing logic for each task is enhanced as follows:
dicedice
prime_factorizationprime\_factorization
palindrome_generationpalindrome\_generation
The ground-truth format expects palindromic character strings (sequences that read the same forward and backward). We remove excess whitespace to address frequent spacing issues in base model responses. This transformation converts spaced responses like “k h g a g h k” to “khgaghk”, preventing failures in string reversibility judgment that occur when spaces interfere with palindrome verification.
advanced_geometryadvanced\_geometry
[…]
power_functionpower\_function
The ground-truth format uses scientific notation. We convert mixed and arithmetic symbol scientific notation to ensure format consistency. The system transforms patterns like $2.36\times 10^{-16}$” or $1.5\times 10^{5}$” to e-notation format (“-2.36e-16”, “1.5e5”), preventing numerically correct but format-incompatible evaluation errors when base models use mixed and arithmetic symbols for scientific notation.
arc_1darc\_1d
[…]
boxnetboxnet
The ground truth format requires dictionary list formatting [{key: value}, ...]. We implement comprehensive cleaning of the JSON format to meet these evaluation requirements. Our processing pipeline includes several steps: rejecting pure numerical responses to prevent non-JSON format interference; removing JSON markdown wrappers that eliminate ‘‘‘json {content} ‘‘‘ markers; converting single dictionaries to single-element dictionary lists (dict $\rightarrow$ [dict]); and filtering illegal elements by removing non-dictionary components from JSON lists. Additionally, we clean nested structure values within individual dictionary entries. For nested lists, we extract the first element as the value ([{key1: [value1, value2, ...]}, ...] $\rightarrow$ [{key1: value1}, ...]). For nested dictionaries, we select matching key values when available ([{key1: {key1: value1, key2: value2, ...}}, ...] $\rightarrow$ [{key1: value1}, ...]) or default to the first element value when keys don’t match ([{key1: {key2: value2, key3: value3}}, ...] $\rightarrow$ [{key1: value2}, ...]). These modifications preserve the model response content to the maximum extent while ensuring compliance with the ground-truth format.
B.4 Entropy Analysis
Setup.
Models.
We evaluate a diverse set of reasoning models to understand entropy characteristics across different training paradigms and parameter scales, as summarized in the table below.
Table 13: Models evaluated in the entropy analysis.
Name Full Model Name Type Parameters
DeepSeek-1.5B DeepSeek-R1-Distill-Qwen-1.5B Base 1.5B
ProRL-1.5B Nemotron-Research-Reasoning-Qwen-1.5B RLVR 1.5B
DeepSeek-7B DeepSeek-R1-Distill-Qwen-7B Base 7B
AceReason-7B AceReason-Nemotron-7B RLVR 7B
Skywork-OR1-7B Skywork-OR1-7B RLVR 7B
DeepSeek-14B DeepSeek-R1-Distill-Qwen-14B Base 14B
AceReason-14B AceReason-Nemotron-14B RLVR 14B
Qwen2.5-32B Qwen2.5-32B Base 32B
DAPO-32B DAPO-Qwen-32B RLVR 32B
Entropy Computation.
For token-level entropy computation, we use teacher forcing to obtain probability estimates. Specifically, after generating the 32 completions with the specified sampling parameters, we feed each generated sequence back to the model and perform a single forward pass to compute the probability distribution over the vocabulary at each token position. Answer-level entropy is computed by first extracting the final answer from each completion using Lighteval (Habib et al., 2023), then calculating the entropy over the distribution of unique answers across the 32 completions. This approach allows us to compute both token-level and answer-level entropy directly from the model’s probability distributions without introducing additional sampling variance.
@@ C.1 Support Preservation: Why RLVR Rarely Discovers New Modes
By construction, we initialize the RLVR policy to the base model as $\pi_{\theta_{0}}(y\mid x)\;=\;q(y\mid x)$. Hence
@@
This manifests as a trade-off between sharpness and diversity: RLVR can improve pass@1 by concentrating mass on known high-reward modes but tends to reduce pass@k performance for larger $k$, where broader coverage is beneficial. By contrast, the base model may occasionally sample correct answers from its long-tail distribution, giving it a statistical edge under high-$k$ evaluations (Yue et al., 2025a; Liu et al., 2025a). This asymptotic upper bound captures a ceiling: no matter how many samples are drawn, the RLVR-trained model cannot exceed the base model’s pass@k in the limit.